OpenStudy (anonymous):

Is X^2 - 16x + 63 prime? What is the factored version?

OpenStudy (anonymous):

what is x equal to ???

OpenStudy (anonymous):

I don't know

OpenStudy (karatechopper):

Set up the parenthesis ( ) ( ) Distribute x^2 (x ) (x ) Now you need to find factors of 63 that will add to be -16. Can you name the factors of 63 for me please?

OpenStudy (anonymous):

umm, 1, 3, 7, 9, 21, 63?

OpenStudy (karatechopper):

Okay. Good don't forget the negative factors as well!! -1, -3,- 7, -9, -21, -63

OpenStudy (anonymous):

(x -21 )(x-3 )

OpenStudy (anonymous):

I am not sure whether that is the correct answer

OpenStudy (anonymous):

does that mean it is prime?

OpenStudy (anonymous):

Yes it is prime.

OpenStudy (anonymous):

No it isn't

hartnn (hartnn):

not actually. from those factors, pick up 2 numbers whose sum is -16 and product = 63

OpenStudy (anonymous):

so its x^2 - 16x +63 = 63?

hartnn (hartnn):

what was that ? 2 numbers from -1, -3,- 7, -9, -21, -63 that add up to -16 are ... ?

OpenStudy (anonymous):

-7, and -9

OpenStudy (anonymous):

( x-7)(x-9) \

OpenStudy (anonymous):

I figured it out all of a sudden

OpenStudy (karatechopper):

Sorry, my wifi cut off..good to see the others stepped in to help..

hartnn (hartnn):

good, so you split -16x as -7x-9x x^2 - 16x +63 = x^2 - 9x-7x +63 factor x from 1st 2 terms and -7 from last two terms, what u get ?

OpenStudy (anonymous):

I figured it out all of a sudden in my mind when I was doing something around the house

OpenStudy (anonymous):

( -7)(-9)=63 -7-9=-16

OpenStudy (anonymous):

now i'm totally lost. Is -16 the factored value? x = -16?

OpenStudy (karatechopper):

x shouldn't equal anything at the moment, you just needed to factor.

OpenStudy (anonymous):

(x-7)(x-9) FOIL First Outer Inner Last Checking stage x^2-9x-7x+63 x^2-16x+63

OpenStudy (anonymous):

x^2 is x squared

OpenStudy (anonymous):

so it is prime?

OpenStudy (anonymous):

It is prime because it can be factored

OpenStudy (anonymous):

thanks!

hartnn (hartnn):

It is NOT prime because it can be factored

OpenStudy (anonymous):

I meant not prime

OpenStudy (anonymous):

I typed in prime

OpenStudy (anonymous):

so what does x^2 - 16x +63 equal?

hartnn (hartnn):

you split -16x as -7x-9x x^2 - 16x +63 = x^2 - 9x-7x +63 factor x from 1st 2 terms and -7 from last two terms, what u get ?

OpenStudy (anonymous):

If you let \(x=1\) doesn't it end up being an even number?